OpenStudy (anonymous):

Let f(x)=x^2-18 and g(x)=6-x. Perform the composition or operation indicated. (fg)(3)

hero (hero):

(fg)(x) = f(x)*g(x) (fg)(3) = (3^2-18)(6-3) = (9 - 18)(3) = (-9)(3) = ?

OpenStudy (anonymous):

-27